# Can It Edit? Evaluating the Ability of Large Language Models to Follow Code Editing Instructions

### 05 Mar 2024 v4

@@ Abstract.
A significant amount of research is focused on developing and evaluating large language models for a variety of code synthesis tasks. These include synthesizing code from natural language instructions, synthesizing tests from code, and synthesizing explanations of code. In contrast, the behavior of instructional code editing with LLMs is understudied. These are tasks in which the model is instructed to update a block of code provided in a prompt. The editing instruction may ask for a feature to added or removed, describe a bug and ask for a fix, ask for a different kind of solution, or many other common code editing tasks.
We introduce a carefully crafted benchmark of code editing tasks and use it to evaluate several cutting edge LLMs. Our evaluation exposes a significant gap between the capabilities of state-of-the-art open and closed models. For example, even GPT-3.5-Turbo is 8.8% better than the best open model at editing code.
We also introduce a new, carefully curated, permissively licensed training set of code edits coupled with natural language instructions. Using this training set, we show that we can fine-tune open Code LLMs to significantly improve their code editing capabilities.
1. Introduction
Instruction Provided to the Model
Edit the C4 class and its methods to represent the C8 group.
Code Diff Between Before and After Segments
@@ 1. Introduction
-class C4(nn.Module): +class C8(nn.Module): -          """Represents the C4 class of group theory, +          """Represents the C8 class of group theory,         where each element represents a discrete rotation."""      def __init__(self):          super().__init__()      def size(self):          """Outputs the size of this group.""" -             return 4 +             return 8      def elements(self):          """Returns all the elements of this group""" -             return torch.tensor([0., np.pi/2, np.pi, 3*np.pi/2]) +             d = np.pi / 4 +             return torch.tensor([0., d, d*2, d*3, d*4, d*5, d*6, d*7])
@@
Figure 1. An abbreviated example of a code editing task from the CanItEdit dataset (Figure 8 presents the full example). The model is tasked with editing the C4 group to represent C8 instead. The model is expected to infer the after code segment from the instruction and the before code segment, as shown in the inferred code diff.
@@
Large language models of code (Code LLMs) are starting to become an essential tool for software engineering practice and research. There has been significant research on synthesizing code from natural language instructions, but comparatively less attention has been given to code editing tasks. However, LLM users expect models to be capable of editing code. For example, the LMsys dataset of in-the-wild conversations with chatbots (Zheng et al., 2023) has 4,188 conversations with code, and 831 (19%) of these involve edits, where the user prompts the model to update generated code based on natural language instructions (Appendix D). In general, code editing encompasses activities like feature addition or removal, bug fixing, and code refactoring (Zhang et al., 2023; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023).
@@
The ability to edit code is also essential for a model to be useful for an AI-focused code editor such as Cursor (Cursor, 2023), Copilot Chat (Copilot, 2023), or ChatGPT Advanced Data Analysis (ADA) (OpenAI, 2023b). Cursor and Copilot Chat facilitate edits with human-written instructions. In contrast, ADA uses both human-written instructions and model-generated reflections (Shinn et al., 2023) to extend and edit code. This approach represents a step towards fully AI-driven code assistance. In both scenarios, instructional code editing is employed, which we define as a function $M(c,I)\rightarrow c^{\prime}$, where $c$ is the original code, $I$ is the instruction, and $c^{\prime}$ is the modified code. An example of this process can be seen in Figure 1, illustrating how the model edits a code segment from a given instruction.
@@
Model-generated reflections and human-written instructions both describe desired code changes. However, they differ in the level of detail: reflections, usually more detailed, are generated by a model with access to the code, offering richer context and potentially a strategic plan for code modifications. In contrast, human-written instructions are typically shorter and less detailed but may express the true user’s intent more clearly. We refer to these as descriptive and lazy instructions, respectively.
@@
In this work, we introduce CanItEdit, a novel dataset comprising 54 hand-crafted instructional code editing problems. These problems, featuring both descriptive and lazy instructions, are coupled with an extensive hidden test suite. Designed to assess a model’s proficiency in handling diverse code editing scenarios, CanItEdit serves as a platform for evaluating state-of-the-art Code LLMs in instructional code editing. Our evaluation focuses on measuring the accuracy of a given model’s ability to write correct code modifications without introducing superfluous code. We conduct comprehensive assessments of closed and open models, revealing significant performance disparities between the leading closed and open models in this domain (Section 5). To help address this gap, we propose a training dataset and methodology for code editing. Our findings demonstrate that fine-tuning open Code LLMs on this dataset can significantly enhance their performance (Section 4).
To summarize, we make the following contributions:
  1. (1)
    We introduce CanItEdit, an extensive and detailed collection of instructional code editing problems, designed to test a model’s ability to edit code under two levels of instruction detail (Section 3).
  2. (2)
    We propose a novel metric, ExcessCode, for assessing code editing models. This metric quantifies the volume of unused code produced by a model when generating a correct solution (Section 5.1).
  3. (3)
    We perform a thorough evaluation of the latest Code LLMs in the context of code editing, providing insights into their current capabilities (Section 5).
  4. (4)
    We present a specially tailored dataset for code editing, along with an effective training methodology, demonstrating significantly enhanced code editing performance through fine-tuning models of three varying sizes (Section 4).
2. Related Work
Instruction-following Language Models
@@ 2. Related Work
Correctly prompting an LLM is crucial for it to perform a desired task. There are multiple methods for instruction tuning LLMs to better adhere to natural language instructions. One method involves employing human annotators to create sample instructions and provide feedback on numerous model outputs (Ouyang et al., 2022; Köpf et al., 2023). However, this method is costly and demands substantial resources. An alternative, cost-effective method is to enable a proficient LLM to self-instruct, generating instructions from a smaller set of human-written seed instructions (Wang et al., 2023). These methods have been applied to generate datasets for instruction-tuning Code LLMs (Chaudhary, 2023; Muennighoff et al., 2023; Luo et al., 2023). Specific to code generation, another strategy to instruction tune an LLM is to use commit messages as instructions (Muennighoff et al., 2023). In this paper, we use commit messages as instructions for code editing. With regards to instruction-tuned models, our results demonstrate that while these models can edit code, they are not as effective as models that are explicitly trained for this task (Section 5).
Code Generation Benchmarks
@@
Several benchmarks exist that test a model’s code generation ability. HumanEval and MBPP are two prominent benchmarks for evaluating Code LLMs in Python programming (Chen et al., 2021; Austin et al., 2021). MultiPL-E expands these benchmarks to 18+ additional programming languages (Cassano et al., 2023b). These benchmarks assess model-generated candidate completions against a series of human-authored unit tests. EvalPlus (Liu et al., 2023) utilizes mutation testing to expand the test suites of the Python benchmarks. All of these benchmarks utilize the pass@k metric, which measures the likelihood of the model generating a completion that passes all of the tests in $k$ tries; we also adopt this metric in our evaluation (Section 5.1). However, these benchmarks are limited to the evaluation of a model’s ability to generate a single function from a natural language description and do not assess code editing capabilities. HumanEvalPack (Muennighoff et al., 2023) is a comprehensive benchmark designed for evaluating Code LLMs across various code generation tasks, such as synthesis, explanation for code understanding, and bug fixing. Specifically, HumanEvalFix, a bug-fixing variant of HumanEvalPack, is extensively used for assessing the models’ capabilities in code refinement (Moon et al., 2023; Muennighoff et al., 2023).
SWE-Bench (Jimenez et al., 2023) evaluates Code LLMs on a broad spectrum of tasks that are performed in the wild by software engineers, and require planning, retrieval, code editing, and more for successful task completion. Our work is more narrowly focused on code editing, and we believe this focus will help guide model development. Another difference with SWE-Bench is that our benchmark is handcrafted, whereas SWE-Bench is based on PRs and issues from popular GitHub repositories. This increases the risk of contamination, particularly with models such as StarCoder, which is trained on several GBs of GitHub issues (Li et al., 2023a).
Code Editing Using Large Language Models
@@
Previous studies on code editing with large language models (LLMs) have predominantly focused on bug fixing (Zhang et al., 2023; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023; Joshi et al., 2023; Wei et al., 2023), a specific subset of code editing, fill-in-the-middle code completion (Bavarian et al., 2022; Fried et al., 2023; Yee and Guha, 2023; Roziere et al., 2023; Guo et al., 2024), an inference strategy that requires specific insert locations, and intrinsic code editing (Li et al., 2023b; Gupta et al., 2023), which involves editing code without a specified instruction, exerting the model’s ability to intrinsically ascertain the desired code changes. Recently, LLMs have progressed in code editing guided by natural language without specific edit locations (Hu et al., 2023; Li et al., 2023a; Muennighoff et al., 2023). However, this advancement lacks benchmark evaluations to effectively measure the models’ code editing skills. Notably, StarCoder (Li et al., 2023a), the first LLM trained on an extensive dataset of commits using the format <before><commit message><after>, has shown enhanced code editing capabilities (Section 5). Before this study, StarCoder’s instructional code editing performance had not been evaluated. The recent introduction of InstructCoder (Hu et al., 2023), a model explicitly trained and evaluated for code editing, marks a significant step towards code editing with LLMs. However, its evaluation involved GPT-4 (OpenAI, 2023a) and human-provided labels on an unreleased dataset, which raises issues regarding reproducibility and comparability in future research and the model has not been publicly released, prohibiting us from evaluating it on our benchmark.
3. The CanItEdit Dataset
Before Code Segment
def hello_world(name):
    return f’{name} says, "Hello World!"’
Lazy Instruction
Make the name fully uppercase.
Descriptive Instruction
The function hello_world should return the string parameter ”name” converted to uppercase concatenated to the string ‘ says, ”Hello World!”’. For example, hello_world(‘the cow’) should return ‘THE COW says, ”Hello World!”’. For another example, hello_world(‘joe’) should return ‘JOE says, ”Hello World!”’.
Reference After Solution
def hello_world(name):
    return f’{name.upper()} says, "Hello World!"’
Hidden Test Suite
hello_world("Bob") == ’BOB says, "Hello World!"’
hello_world("") == ’ says, "Hello World!"’
hello_world("Joe") == ’JOE says, "Hello World!"’
Edit Kind: Revise
Topic: Language Processing
Figure 2. The hello_world problem from CanItEdit. This problem is the easiest in the dataset, and is intended to be used as a sanity check.
This section presents CanItEdit, a dataset of Python code editing problems with natural language instructions, hand-written and cross-validated by experienced computer science experts for evaluating LLMs’ code editing capabilities.
3.1. Problem Construction
CanItEdit features 54 Python code editing problems, each comprising a ‘before’ and an ‘after’ code segment, two types of natural language instructions (descriptive and lazy), and a hidden test suite. The task for models is to transform the ‘before’ code segment into the ‘after’ segment based on either instruction, aiming to pass the hidden tests. Inspired by HumanEval’s methodology (Austin et al., 2021), we hand-wrote the problems, avoiding public sources like GitHub to reduce pre-training exposure. We also verified that the instructions are unique to this dataset and not part of our fine-tuning data (section 4). Problems range from simple function edits to complex, multi-class challenges, covering data structures, algorithms, mathematics, language processing, and game programming. Some require popular external Python libraries like NumPy, Pandas, PyTorch, and Z3. Dataset statistics and example problems are detailed in Table 1 and Appendix B, respectively.
The ‘before’ code segments in CanItEdit represent various starting states, ranging from functional programs needing additional features to those with bugs or incomplete implementations requiring fixes or optimizations. Conversely, the ‘after’ segments illustrate the correct solutions that fulfill the task requirements and clear the test suite.
We categorize code editing tasks into two types: evolve and revise. Evolve tasks involve adding or removing major features like new methods or classes. In contrast, Revise tasks focus on modifying existing functionalities, including bug fixing, logic changes, or refactoring, such as transitioning from imperative to object-oriented programming, as exemplified in Figure 7. The distinction between these categories is based on the edit’s primary objective, though some tasks may exhibit characteristics of both.
The dataset’s dual natural language instructions test model efficiency in two scenarios: 1) Descriptive: Detailed instructions replicate situations where users provide specific specifications or another model outlines a plan, similar to Reflexion prompting (Shinn et al., 2023; Fan et al., 2023; Phung et al., 2023). 2) Lazy: Informal instructions resemble typical user queries for LLMs in code generation (Babe et al., 2023).
In both, the model must generate code that meets the instruction and passes the hidden tests. Descriptive instructions offer detailed guidance, including function names and input-output examples, while lazy instructions provide minimal information, requiring the model to infer user intent and rely more on the ‘before’ segment. Both instructions should lead to an equivalent ‘after’ segment. For instance, in the hello_world problem (Figure 2), the descriptive instruction is comprehensive, whereas the lazy instruction is brief, pushing the model to deduce user intent. Further discussions on human-written and Reflexion-generated instructions are in Appendix D.
3.2. Test Suites
For our test suites, we ensure three essential properties:
  1. (1)
    Completeness: Each suite comprehensively covers many inputs and edge cases. This includes numerous test cases per problem, targeting edge and corner cases, with $100\%$ code coverage verified using Coverage.py (Batchelder and Contributors to Coverage.py, [n. d.]). It is worth noting that code coverage is not as robust as mutation testing, which is employed by EvalPlus (Liu et al., 2023).
  2. (2)
    Correctness: The suites are bug-free, passing all tests with the ‘after’ code while failing at least one with the ‘before’ code.
  3. (3)
    Concealment: Test suites are hidden from models during training and inference, achieved by excluding them from the training dataset and not presenting them during model evaluation.
We hand-crafted test suites for each problem, incorporating diverse testing methods ranging from simple unit tests to complex property-based testing, mocking, fuzzing, and integration tests. For instance, one of our benchmark problems involves implementing a strategy for a Tic-Tac-Toe game that outperforms a baseline strategy (Figure 9). The lazy instruction for this problem is: Create a strategy ‘GoodStrategy’, that beats ‘CornerStrategy’. Do not modify the ‘Game’ class. To test this, the suite includes unit tests for both the ‘Game’ and ‘CornerStrategy’ classes, along with integration tests that evaluate the entire program, ensuring that ‘GoodStrategy’ wins over ‘CornerStrategy’. Additionally, we use Python’s inspect module to check if the ‘Game’ class remains unmodified, adhering to the problem’s constraints.
@@ 3.2. Test Suites
Total Problems (Revise/Evolve) 54 (32/22)
@@
DS & Algorithms 22
Language Processing 15
Mathematics 9
Game Programming 8
External Library Usage
NumPy (6), PyTorch (2), Pandas (1), Z3 (2)
Code Segment Mean Std. Dev.
Mean Lines (Before/After) 44.7/53.6 38.5/40.7
Levenshtein Distance 361.2 387.1
Combined Mean Lines 98.3 78.4
Combined Mean Tokens 888.6 702.6
@@
Instruction Mean Std. Dev.
Mean Tokens (Descriptive/Lazy) 89.0/40.3 57.6/35.6
Table 1. Dataset statistics for CanItEdit.
4. Fine-tuning
This section outlines our methodology for fine-tuning a Code LLM specifically for code editing tasks. We fine-tune models based on the DeepSeek-Coder-Base family of Code LLMs (Guo et al., 2024), which are variants of CodeLlama (Roziere et al., 2023) trained from scratch on 2T tokens comprised of 87% permissively licensed code from GitHub and 13% natural language, using the same filtering rules as StarCoder’s data collection (Li et al., 2023a). At the time of writing, these models are the top-performing, open-access foundational Code LLMs, excelling in various code generation benchmarks. Furthermore, they are distributed under a permissive open-source license, allowing free use and modification for research and commercial purposes. We selected these base models because they exhibit robust performance on CanItEdit, even without being specifically trained for this or any other instructional tasks, thus highlighting their exceptional generalization capabilities (Section 5).

### 20 Mar 2024 v5

@@ Abstract
A significant amount of research is focused on developing and evaluating large language models for a variety of code synthesis tasks. These include synthesizing code from natural language, synthesizing tests from code, and synthesizing explanations of code. In contrast, the behavior of instructional code editing with LLMs is understudied. These are tasks in which the model is provided a block of code and an instruction to modify the code. The editing instruction may ask for a feature to be added or removed, describe a bug and ask for a fix, or ask for a different kind of solution. We introduce a carefully crafted benchmark of code editing tasks and use it to evaluate several cutting edge LLMs. Our evaluation exposes a significant gap between the capabilities of state-of-the-art open and closed models. For example, even GPT-3.5-Turbo is better than the best open model at code editing tasks. We also introduce a new, carefully curated, permissively licensed training dataset of code editing tasks coupled with natural language instructions. Using this training dataset, we show that we can fine-tune open Code LLMs to significantly improve their code editing capabilities, closing the gap between open and closed models. All code, data, and models are available at https://github.com/nuprl/CanItEdit.
Federico Cassano Luisa Li Akul Sethi
Northeastern University Northeastern University Northeastern University
cassano.f@northeastern.edu
Noah Shinn Abby Brennan-Jones Jacob Ginesin
Northeastern University Wellesley College Northeastern University
Edward Berman George Chakhnashvili Anton Lozhkov
Northeastern University Northeastern University HuggingFace
Carolyn Jane Anderson Arjun Guha
Wellesley College Northeastern University
a.guha@northeastern.edu
1 Introduction
Instruction: Edit the C4 class and its methods to represent the C8 group instead.
@@ 1 Introduction
-class C4(nn.Module): +class C8(nn.Module): -     """Represents the cyclic group C4, +     """Represents the cyclic group C8,         where each element represents a discrete rotation."""      def __init__(self):          super().__init__()      def size(self):          """Outputs the size of this group.""" -           return 4 +           return 8      def elements(self):          """Returns all the elements of this group""" -           return torch.tensor([0., np.pi/2, np.pi, 3*np.pi/2]) +           d = np.pi / 4 +           return torch.tensor([0., d, d*2, d*3, d*4, d*5, d*6, d*7])
@@
Figure 1: An abbreviated example of a code editing task from the CanItEdit dataset (Figure 9 presents the full example). The model is tasked with editing the C4 group to represent C8 instead. The model is expected to infer the after code segment from the instruction and the before code segment, as shown in the inferred code diff.
@@
Large language models of code (Code LLMs) are becoming an essential tool for software engineering practice and research. There has been significant research on synthesizing code from natural language instructions, but comparatively less attention has been given to code editing tasks. However, LLM users expect models to be capable of editing code. For example, the LMsys dataset of in-the-wild conversations with chatbots (Zheng et al., 2023) has 4,188 conversations containing code, and 831 (19%) of these involve code editing, where the user prompts the model to update code based on natural language instructions (Appendix E). In general, code editing with an LLM encompasses activities like feature addition or removal, bug fixing, and code refactoring (Zhang et al., 2023; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023).
@@
The ability to edit code is also essential for a model to be useful for an AI-focused code editor such as Cursor (Cursor, 2023), Copilot Chat (Copilot, 2023), or ChatGPT Advanced Data Analysis (ADA) (OpenAI, 2023a). Cursor and Copilot Chat facilitate edits with human-written instructions. In contrast, ADA uses both human-written instructions and model-generated reflections (Shinn et al., 2023; Fan et al., 2023; Phung et al., 2023) to extend and edit code. This approach represents a step towards AI-driven code assistance. In both scenarios, instructional code editing is employed, which we define as a function $M(c,I)\rightarrow c^{\prime}$, where $c$ is the original code, $I$ is the instruction, and $c^{\prime}$ is the modified code. Figure 1 illustrates this process, showing how the model edits a code segment from a given instruction.
@@
Model-generated reflections and human-written instructions both describe desired code changes. However, they differ in the level of detail: reflections, usually more detailed, are generated by a model with access to the code, offering richer context and potentially a strategic plan for code modifications. In contrast, human-written instructions are typically shorter and less detailed but may express the true user’s intent more clearly. We refer to these as descriptive and lazy instructions, respectively. We thoroughly analyze examples of such instructions in Appendix E.
@@
In this work, we introduce CanItEdit, a novel dataset comprising 105 hand-crafted instructional code editing problems, featuring both descriptive and lazy instructions and an extensive hidden test suite. Designed to assess a model’s proficiency in handling diverse code editing scenarios, CanItEdit serves as a benchmark for evaluating state-of-the-art Code LLMs in instructional code editing. Our evaluation focuses on measuring the accuracy of a given model’s ability to write correct code modifications without introducing unnecessary code. We conduct comprehensive assessments of closed and open models, revealing significant performance disparities between the leading closed and open models (§ 5). To help address this gap, we propose a training dataset and methodology for code editing. Our findings demonstrate that fine-tuning open Code LLMs on this dataset can significantly enhance code editing performance (§ 4).
To summarize, we make four main contributions: (1) We introduce CanItEdit, a dataset of instructional code editing problems, designed to evaluate the code editing capabilities of large language models (§ 3). (2) We propose a novel metric, ExcessCode, for quantifying the typical volume of unused code produced by a model when generating the correct code edits (§ 5.1). (3) We perform a thorough evaluation of the latest Code LLMs in the context of code editing, providing insights into their current capabilities (§ 5). (4) Finally, we present a specially tailored training dataset for code editing, along with an effective training methodology, demonstrating significantly enhanced code editing performance through fine-tuning models of varying sizes (§ 4).
2 Related Work
@@
Instruction-following Language Models. Correctly prompting an LLM is crucial for it to perform a desired task. There are multiple methods for instruction tuning LLMs to better adhere to natural language instructions. One method involves employing human annotators to create sample instructions and provide feedback on numerous model outputs (Ouyang et al., 2022; Köpf et al., 2023). This method is costly and demands substantial resources. An alternative, cost-effective method is to use an LLM to self-instruct, generating instructions from a smaller set of human-written seed instructions (Wang et al., 2023). These methods have been applied to generate datasets for instruction-tuning Code LLMs (Chaudhary, 2023; Luo et al., 2023). Specific to code generation, another strategy to instruction-tune an LLM is to use commit messages as instructions (Muennighoff et al., 2023). In this paper, we use commit messages as instructions for code editing. Our results demonstrate that while instruction-tuned models can edit code, they are not as effective as models that we explicitly train for this task (§ 5).
@@
Code Generation Benchmarks. Several benchmarks exist that test a model’s code generation ability. HumanEval and MBPP are two prominent benchmarks for evaluating LLMs in Python programming (Chen et al., 2021; Austin et al., 2021). MultiPL-E expands these benchmarks to 18+ additional programming languages (Cassano et al., 2023b). These benchmarks assess model-generated candidate completions against a series of human-authored unit tests. EvalPlus (Liu et al., 2023) utilizes mutation testing to expand the test suites of the Python benchmarks. All of these benchmarks utilize the pass@k metric, which measures the likelihood of the model generating a completion that passes all of the tests in $k$ tries; we also adopt this metric in our evaluation (§ 5.1). However, these benchmarks are limited to the evaluation of a model’s ability to generate a single function from a natural language description and do not assess code editing capabilities. HumanEvalPack (Muennighoff et al., 2023) is a benchmark designed for evaluating LLMs across various single-function code generation tasks, such as synthesis, code explanation, and bug fixing. Specifically, HumanEvalFix, a bug-fixing variant of HumanEvalPack, is extensively used for assessing the models’ capabilities in code refinement (Moon et al., 2023; Muennighoff et al., 2023). However, the instruction is fixed for every problem. SWE-Bench (Jimenez et al., 2023) evaluates LLMs across varied programming tasks including planning, retrieval, and code editing. Our work concentrates specifically on code editing tasks, aiming to more precisely guide model development. Unlike SWE-Bench, which sources its problems from GitHub PRs and issues, our benchmark is handcrafted, reducing contamination risks as seen with models like StarCoder and StarCoder2, which are trained extensively on GitHub data (Li et al., 2023b; Lozhkov et al., 2024).
@@
Code Editing Using Large Language Models. Previous studies on code editing with LLMs have predominantly focused on bug fixing (Zhang et al., 2023; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023; Joshi et al., 2023; Wei et al., 2023), a specific subset of code editing; fill-in-the-middle code completion (Bavarian et al., 2022; Fried et al., 2023; Yee & Guha, 2023; Roziere et al., 2023; Guo et al., 2024), an inference strategy that requires specific insert locations; and intrinsic code editing (Li et al., 2023a; Gupta et al., 2023), which involves editing code without a specified instruction, exerting the model’s ability to intrinsically ascertain the desired code changes. Recently, LLMs have progressed in code editing guided by natural language without specific edit locations (Hu et al., 2023; Li et al., 2023b; Muennighoff et al., 2023). However, this advancement lacks benchmark evaluations to effectively measure the models’ code editing skills. Notably, StarCoder (Li et al., 2023b), the first LLM trained on an extensive dataset of commits using the format <before><commit message><after>, we have shown enhanced code editing capabilities (§ 5). Before this study, StarCoder’s practical code editing performance had not been assessed. StarCoder2 has replaced commits with pull requests and issues, which typically include more natural language (Lozhkov et al., 2024). The recent introduction of InstructCoder (Hu et al., 2023), a model explicitly trained and evaluated for code editing, marks a significant step towards code editing with LLMs. However, its evaluation involved GPT-4-generated (OpenAI, 2023b) and human-provided labels, which raises issues regarding reproducibility and comparability in future research. Moreover, the model has not been publicly released, prohibiting us from evaluating it on our benchmark.
3 The CanItEdit Dataset
@@
Total Tasks 105 (35/35/35)
Total Problems 210 (70/70/70)
@@
Data Structures & Algorithms 39
Language Processing 21
Mathematics 25
Data Science 10
Miscellaneous 10
Problems With Library Usage 22
Code Segment Mean $\pm$ Std. Dev.
[…]
Levenshtein Distance 302.1 $\pm$ 339.6
Combined Mean Lines 92.3 $\pm$ 69.9
Combined Mean Tokens 865.3 $\pm$ 639.7
@@
Instruction Mean $\pm$ Std. Dev.
[…]
Table 1: Dataset statistics for CanItEdit.
Benchmark Overview
CanItEdit is a dataset comprising 105 meticulously constructed Python code editing challenges. Each problem includes the input code segment (before), the expected code segment (after), the two types of natural language instructions (descriptive and lazy), and a hidden test suite. These challenges span a broad spectrum of computer science domains, such as data structures, algorithms, mathematics, language processing, and game programming, requiring knowledge of popular external Python libraries like NumPy, Pandas, PyTorch, and others. Table 1 presents general dataset statistics for CanItEdit.
Following Swanson (1976) and follow-up work (Levin & Yehudai, 2017), we classify code editing tasks into three distinct categories based on their primary goal: a corrective edit fixes errors, a perfective edit enhances existing features, and an adaptive edit meets new requirements. We have 35 problems per category for an even distribution across the different types of code changes. The dual instruction formats test the models’ ability to execute tasks based on the provided context: descriptive instructions provide comprehensive details for explicit guidance, while lazy instructions offer minimal direction, challenging the model to infer the required actions. Both instructions should lead to an equivalent after segment. Descriptive instructions serve to replicate situations where users provide specific specifications or another model outlines a plan. In contrast, lazy instructions resemble typical user queries for LLMs in code generation. As each problem has two distinct instructions, the dataset effectively contains 210 problems. We showcase examples from CanItEdit in Appendix C.
Dataset Creation
To manually construct CanItEdit, we assembled a team of eight experienced Python programmers, each with different domain expertise, and appointed one as the lead. Our objective was to fill each change category with 35 problems, with 105 problems total. Before starting, we provided the team with verbal and written guidance, a standard template, and an example problem. They were instructed to begin by writing a brief description of the problem and the applied changes for review and refinement by the lead. Next, they were tasked to write the ’before’ code segment and hidden test suite, followed by the ’after’ code segment, along with the instructions. The lead initially reviewed all problems in development and they were additionally reviewed by the entire team in weekly meetings. Upon completion of a problem, the lead generated sample completions to ensure that the failures and successes were reasonable and consistent with the problem’s intent.
The team also dedicated significant effort to developing comprehensive test suites for each problem, which incorporated a variety of testing techniques such as unit tests, property-based testing, mocking, fuzzing, and integration tests. These suites were designed to rigorously evaluate whether the ’after’ segment met the problem requirements while ensuring the ’before’ code did not. To confirm the completeness and correctness of the test suites, we created an automated verification pipeline that ensured $100\%$ line coverage and that the suite passed all tests with the ‘after’ code while failing at least one with the ‘before’ code. The team also manually reviewed the tests to ensure correctness and completeness.
4 Fine-Tuning
We describe our approach to fine-tuning Code LLMs for code editing tasks, focusing on the DeepSeekCoder-Base family (Guo et al., 2024), a variant of CodeLlama (Roziere et al., 2023) trained on 2 trillion tokens of GitHub code and natural language, using StarCoder’s filtering rules (Li et al., 2023b). These models, top-performing in code generation and open-access under a permissive license, show robust performance on CanItEdit without specific training for instructional tasks (§ 5).